OpenStudy (11study11):

Part A: Sam rented a boat at $225 for 2 days. If he rents the same boat for 5 days, he has to pay a total rent of $480. Write an equation in the standard form to represent the total rent (y) that Sam has to pay for renting the boat for x days. Part B: Write the equation obtained in Part A using function notation. Part C: Describe the steps to graph the equation obtained above on the coordinate axes. Mention the labels on the axes and the intervals

OpenStudy (11study11):

@TheSmartOne

OpenStudy (11study11):

@jim_thompson5910

OpenStudy (11study11):

@zepdrix

jimthompson5910 (jim_thompson5910):

@11study11 how far did you get?

jimthompson5910 (jim_thompson5910):

or were you not able to get started?

OpenStudy (11study11):

Part A: y = 85x + 55 Part B: f(x)= 85x + 55 @jim_thompson5910

jimthompson5910 (jim_thompson5910):

ok make sure you show your steps to your teacher. You have the correct result for part A

jimthompson5910 (jim_thompson5910):

also for part B as well

jimthompson5910 (jim_thompson5910):

so you're stuck on how to graph?

OpenStudy (11study11):

yes

jimthompson5910 (jim_thompson5910):

are you able to identify the slope and y intercept?

OpenStudy (11study11):

slope= 85 y intercept=55

jimthompson5910 (jim_thompson5910):

good, so your first point will be (0,55) basically at 55 on the y axis

jimthompson5910 (jim_thompson5910):

the next point will be found by going up 85 units, then over to the right 1 to land on the point (1,140) Notice this is the same if you plugged in x = 1 into y = 85x+55 y = 85*x + 55 y = 85*1 + 55 y = 140

jimthompson5910 (jim_thompson5910):

Plot the two points (0,55) and (1,140). Then draw a straight line through them

OpenStudy (11study11):

ok

OpenStudy (11study11):

now what

jimthompson5910 (jim_thompson5910):

once you draw the line through those two points, you're done with the graph

jimthompson5910 (jim_thompson5910):

don't forget to `Mention the labels on the axes and the intervals`

jimthompson5910 (jim_thompson5910):

I don't think you actually have to graph. It just says to describe how to graph. It seems like your teacher doesn't want the actual graph. However, it's still good practice.

OpenStudy (11study11):

how would i write that?

jimthompson5910 (jim_thompson5910):

Like I wrote out, but in your own words of course.

OpenStudy (11study11):

for part b how would i explain that

jimthompson5910 (jim_thompson5910):

you would simply replace the 'y' with f(x). So there's not much to part B. Much of the work is done in part A

OpenStudy (11study11):

ok thanks

OpenStudy (11study11):

Part A: 225 = 2a + b 480 = 5a + b y = 85x + 55 Part B: f(x)= 85x + 55 Part C: y = 85x + 55 y = 85*1 + 55 y = 140 Plot the two points (0,55) and (1,140).

OpenStudy (11study11):

is it all correct?

jimthompson5910 (jim_thompson5910):

It's not clear how you jumped from this ``` 225 = 2a + b 480 = 5a + b ``` to this `y = 85x + 55`

OpenStudy (11study11):

hmm don't know what I did there.

jimthompson5910 (jim_thompson5910):

how did you get 85? did you use the slope formula?

OpenStudy (11study11):

yup

jimthompson5910 (jim_thompson5910):

ok so show that to your teacher instead of ``` 225 = 2a + b 480 = 5a + b ```

OpenStudy (11study11):

wait..

OpenStudy (11study11):

I'm confused.

jimthompson5910 (jim_thompson5910):

use the slope formula m = (y2-y1)/(x2-x1)

jimthompson5910 (jim_thompson5910):

(x1,y1) = (2,225) (x2,y2) = (5,480) m = (y2-y1)/(x2-x1) m = (480-225)/(5-2) m = (255)/(3) m = 85 which is what I'm assuming what you did but didn't write it in this exact form

OpenStudy (11study11):

OK Thanks!!!

jimthompson5910 (jim_thompson5910):

No problem. Also, show how you got (0,55). So you can mention it's the y intercept or plug in x = 0 like you plugged in x = 1.

OpenStudy (11study11):

ok